Two cards are drawn from a standard deck of cards with replacement.
What is the probability that the first is a King and the second is not a King?

Answer by nerdybill(7384)

Two cards are drawn from a standard deck of cards with replacement.
What is the probability that the first is a King and the second is not a King?
.
Probability of first card:
Number of Kings in a deck: 4
Total number of cards in a deck: 52
probability: 4/52 = 1/13
.
Probability of second card:
Number of cards not a King: 52-4 = 48
Total number of cards in a deck: 52
probability: 48/52 = 12/13
.
Probability of event:
1/13 * 12/13 = 12/169 = .071 =7.1%